given that y1(x)=x1 is 1 solution of differential equation 2 x2 y,, + 3 x y, - y =0 and x>0. find a second end independent solution of the equation


1
Expert's answer
2021-04-28T11:05:18-0400


Look for a solution in the form 


"y=y_1z=\frac{z}{x}",


then "y'=\frac{z'}{x}-\frac{z}{x^2}" 


and "y''=z''\left(\frac{1}{x}\right)+2z'\left(\frac{1}{x}\right)'+z\left(\frac{1}{x}\right)''=\frac{z''}{x}-\frac{2z'}{x^2}+\frac{2z}{x^3}"

Then 


"0=2x^2\left(\frac{z''}{x}-\frac{2z'}{x^2}+\frac{2z}{x^3}\right)+3x\left(\frac{z'}{x}-\frac{z}{x^2}\right)-\frac{z}{x}=2z''x-z'"

Let "z'=u", then "2u'x-u=0" or "\frac{2u'ux-u^2x'}{x^2}=0".



Since "2u'u=(u^2)'", we have "0=\frac{2u'ux-u^2x'}{x^2}=\frac{(u^2)'x-u^2x'}{x^2}=\left(\frac{u^2}{x}\right)'", so "\frac{u^2}{x}=C"

Then "u=\sqrt{Cx}=\sqrt{|C|}\sqrt{x}" if "C\ge 0",


or "u=\sqrt{Cx}=\sqrt{|C|}\sqrt{-x}" if "C\le 0".


Denote "\sqrt{|C|}" by "D", then "u=D\sqrt{x}" or "u=D\sqrt{-x}"



Replace "u" by "z'""z'=D\sqrt{x}" or "z'=D\sqrt{-x}".




Then 

 "z=\frac{2}{3}D\sqrt{x^3}+A" or "z=-\frac{2}{3}D\sqrt{-x^3}+A". In every case "z=B\sqrt{|x|^3}+A".



Since "y=\frac{z}{x}", we have "y=Bs(x)\sqrt{|x|}+\frac{A}{x}=E\sqrt{|x|}+\frac{A}{x}",



where "s(x)=\begin{cases}
1,&\text{if $x>0$}\\
0,&\text{if $x=0$}\\
-1,&\text{if $x<0$}
\end{cases}"



So "\sqrt{|x|}" is the second solution.

Answer: "\sqrt{|x|}"